### v2

# On dissipation operators of Quantum Optics

Abstract
We consider dissipation operators used in Quantum Optics for the description of quantum spontaneous emission in the context of damped driven Jaynes–Cummings equations. The equations describe quantised one-mode Maxwell field coupled to a two-level molecule. Our main result is the symmetry and nonpositivity of basic dissipation operator of Quantum Optics.
MSC classification: 81V80, 81S05, 81S08 37K06, 78A40, 78A60.
Keywords: Quantum Optics; laser; Jaynes–Cummings model; Hamiltonian; density operator; pumping; dissipation; Hilbert–Schmidt operator; trace.
1 Jaynes–Cummings equation
The Jaynes–Cummings equation (JCE) is basic model of Quantum Optics. The model without damping and pumping was introduced in 1963 by Jaynes and Cummings [12] (the survey can be found in [16]). The damped driven Jaynes–Cummings equations read as [21, (5.107)]:
$\dot{\rho}(t)={\cal A}(t)\rho(t):=-i[H(t),\rho(t)]+\gamma D\rho(t),\qquad t\geq 0.$ (1.1)
[…]
$a|n\rangle=\sqrt{n}|n-1\rangle,\qquad a^{\dagger}|n\rangle=\sqrt{n+1}|n+1\rangle,\qquad[a,a^{\dagger}]=1.$ (1.2)
The Hamiltonian $H(t)$ is the sum
$H(t)\!=\!H_{0}\!+\!pV(t),\,\,\,H_{0}\!=\!H_{F}\!+\!H_{A},\,\,\,H_{F}\!=\!\omega_{c}a^{\dagger}a,\,\,\,H_{A}\!=\!\frac{1}{2}\omega_{a}\sigma_{3},\,\,\,V(t)\!=\!\sigma_{1}(a\!+\!a^{\dagger})\!+\!A^{e}(t).$ (1.3)
[…]
2 Dissipation operators and main results
The dissipation operator $D$ in (1.1) describes the quantum spontaneous emission which provides an energy loss to balance the pumping. The dissipation operators must satisfy the following conditions.
We will consider the dissipation operators [21, (5.107)] and [5, (5.6.20)]:
$D\rho=a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a+a^{\dagger}\rho a-\frac{1}{2}aa^{\dagger}\rho-\frac{1}{2}\rho aa^{\dagger},\qquad{\Delta}\rho=a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a.$ (2.1)
Definition 2.1.
[…]
ii) $X_{0}$ is the space of finite linear combinations of the vectors $|n,s_{\pm}\rangle$.
iii) ${\cal D}\subset{\rm HS}$ is the subspace of finite rank Hermitian operators
$\rho=\sum_{n,n^{\prime}}\,\sum_{s,s^{\prime}=s_{\pm}}\rho_{n,n^{\prime}}|n,s\rangle\otimes\langle n^{\prime},s^{\prime}|.$ (2.2)
Every density operator $\rho\in{\rm HS}$ is defined uniquely by its matrix entries
$\rho_{n,n^{\prime}}=\langle n|\rho|n^{\prime}\rangle,$ (2.3)
which are Hermitian $2\times 2$ matrices. The Hilbert–Schmidt norm, corresponding to the inner product (2.5), can be written as
$\|\rho\|_{\rm HS}^{2}=\mathop{\rm tr\,}\nolimits[\rho^{2}]=\sum_{n,n^{\prime}=0}^{\infty}\mathop{\rm tr\,}\nolimits\rho^{2}_{n,n^{\prime}}<\infty.$ (2.4)
Definition $2.$
$\rm HS$} is the Hilbert space of Hermitian Hilbert–Schmidt operators with the inner product [19, Thm VI.22 (c)]
$\langle\rho_{1},\rho_{2}\rangle_{\rm HS}=\mathop{\rm tr\,}\nolimits[\rho_{1}\rho_{2}].$ (2.5)
The operator $D$ is a modification of the version $\Delta$}. The version $\Delta$} has been introduced in [1][3], and used in [5, (5.6.20)] and [6, (18)], [17], [20, (4)]. The version is symmetric in the following two aspects: for $\rho\in{\cal D}$, ${\Delta}\rho$ is the selfadjoint operator with zero trace that is necessary for the dynamics of $\rho(t)$ in selfadjoint operators with constant trace. The modification $D$ restores the third symmetry with respect to the interchange of $a$ and $a^{\dagger}$.
The main result of present paper is the following theorem.
Theorem 2.3.
i) The dissipation operator $D$ is defined on the domain ${\cal D}\subset{\rm HS}$ and it is nonpositive and symmetric:
$\langle\rho,D\rho\rangle_{\rm HS}\leq 0,\qquad\langle\rho,D\rho\rangle_{\rm HS}=\langle D\rho,\rho\rangle_{\rm HS},\qquad\rho\in{\cal D}.$ (2.6)
ii) The operator ${\Delta}:{\cal D}\to{\cal D}$ is neither symmetric nor nonpositive.
iii) The operator $D:{\cal D}\to{\cal D}$ is injective.
Corollary 2.4.
The dissipation operator $D$ admits the selfadjoint Friedrichs extension.
Let us comment on related previous results. The dissipation operators $\Delta$} and $D$ were introduced analysing quantum spontaneous emission in [1][3] and [5, 6, 17, 20, 21].
In [14] and [15], we have demonstrated the role of classical dissipation in adiabatic single-frequency asymptotics for solutions to the semiclassical Maxwell–Bloch equations.
3 Nonpositivity and symmetry
Here we prove Theorem 2.3 i) and ii).
i) First, let us prove the nonpositivity of $D$. For $\rho\in{\cal D}$,
$\langle\rho,D\rho\rangle_{\rm HS}$ $=$ $\mathop{\rm tr\,}\nolimits\Big{(}\rho D\rho\Big{)}=\mathop{\rm tr\,}\nolimits\Big{(}\rho\big{[}a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho(t)-\frac{1}{2}\rho a^{\dagger}a+a^{\dagger}\rho a-\frac{1}{2}aa^{\dagger}\rho-\frac{1}{2}\rho aa^{\dagger}\big{]}\Big{)}$ (3.1)
$=$ $\mathop{\rm tr\,}\nolimits\Big{(}\rho a\rho a^{\dagger}-\rho a^{\dagger}a\rho+\rho a^{\dagger}\rho a-\rho aa^{\dagger}\rho\Big{)}=\mathop{\rm tr\,}\nolimits\Big{(}2\rho a\rho a^{\dagger}-a^{\dagger}a\rho^{2}-aa^{\dagger}\rho^{2}\Big{)}.$
Now we use the fact that $\rho$ is a finite rank Hermitian operator (2.2). Then all operators $\rho a\rho a^{\dagger}$, $a^{\dagger}a\rho^{2}$ and $aa^{\dagger}\rho^{2}$ have only finite number of nonzero entries (2.3), so their traces are well defined. Moreover, $\rho$ admits a finite spectral resolution $\rho=(\rho_{i}\delta_{ij})$ in the orthonormal basis of its eigenvectors from $X_{0}$:
$\rho=\sum_{i=1}^{\infty}\rho_{i}\,e_{i}\!\otimes\!e_{i},\qquad e_{i}\in X_{0},$ (3.2)
where only finite number of the eigenvalues $\rho_{i}$ are nonzero. In this basis, the entries $a_{jk}=\langle e_{j},ae_{k}\rangle$ and $a^{\dagger}_{kl}=\langle e_{k},a^{\dagger}e_{l}\rangle$ of the operators $a$ and $a^{\dagger}$ are well defined. Hence, with summation in repeated indices,
$\langle\rho,D\rho\rangle_{\rm HS}$ $=$ $\Big{(}2\rho_{i}\delta_{ij}a_{jk}\rho_{k}\delta_{kl}a^{\dagger}_{li}-a^{\dagger}_{kl}a_{lj}\rho_{j}^{2}(t)\delta_{jk}-a_{jk}a^{\dagger}_{kl}\rho_{l}^{2}(t)\delta_{lj}\Big{)}$ (3.3)
$=$ $\Big{(}2\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{kl}a_{lk}\rho_{k}^{2}-a_{jk}a^{\dagger}_{kj}\rho_{j}^{2}\Big{)}$
$=$ $\Big{(}2\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{ki}a_{ik}\rho_{k}^{2}-a_{ik}a^{\dagger}_{ki}\rho_{i}^{2}\Big{)}$
$=$ $a_{ik}a^{\dagger}_{ki}[2\rho_{i}\rho_{k}-\rho_{k}^{2}-\rho_{i}^{2}]=-|a_{ik}|^{2}[\rho_{i}-\rho_{k}]^{2}\leq 0$
since $a^{\dagger}_{ji}=\overline{a}_{ij}$. Hence, the nonpositivity of $D$ is proved.
Second, let us check the symmetry of $D$:
$\mathop{\rm tr\,}\nolimits[\rho_{1}D\rho_{2}]=\mathop{\rm tr\,}\nolimits[D\rho_{1})^{{\dagger}}\rho_{2}],\qquad\rho_{1},\rho_{2}\in{\cal D}.$ (3.4)
We have
$\mathop{\rm tr\,}\nolimits[\rho_{1}D\rho_{2}]$ $=$ $\mathop{\rm tr\,}\nolimits[\rho_{1}(a\rho_{2}a^{\dagger}-\frac{1}{2}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{2}a^{{\dagger}}a+a^{{\dagger}}\rho_{2}a-\frac{1}{2}aa^{{\dagger}}\rho_{2}-\frac{1}{2}\rho_{2}aa^{{\dagger}})]$
$=$ $\mathop{\rm tr\,}\nolimits[\rho_{1}a\rho_{2}a^{\dagger}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{1}\rho_{2}a^{{\dagger}}a+\rho_{1}a^{{\dagger}}\rho_{2}a-\frac{1}{2}\rho_{1}aa^{{\dagger}}\rho_{2}-\frac{1}{2}\rho_{1}\rho_{2}aa^{{\dagger}}]$
$=$ $\mathop{\rm tr\,}\nolimits[a^{\dagger}\rho_{1}a\rho_{2}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}a^{{\dagger}}a\rho_{1}\rho_{2}+a\rho_{1}a^{{\dagger}}\rho_{2}-\frac{1}{2}\rho_{1}aa^{{\dagger}}\rho_{2}-\frac{1}{2}aa^{{\dagger}}\rho_{1}\rho_{2}]$
$=$ $\mathop{\rm tr\,}\nolimits[(a^{\dagger}\rho_{1}a-\frac{1}{2}\rho_{1}a^{{\dagger}}a-\frac{1}{2}a^{{\dagger}}a\rho_{1}+a\rho_{1}a^{{\dagger}}-\frac{1}{2}\rho_{1}aa^{{\dagger}}-\frac{1}{2}aa^{{\dagger}}\rho_{1})\rho_{2}]$
$=$ $\mathop{\rm tr\,}\nolimits[(a^{\dagger}\rho_{1}a-\frac{1}{2}a^{{\dagger}}a\rho_{1}-\frac{1}{2}\rho_{1}a^{{\dagger}}a+a\rho_{1}a^{{\dagger}}-\frac{1}{2}aa^{{\dagger}}\rho_{1}-\frac{1}{2}\rho_{1}aa^{{\dagger}})^{{\dagger}}\rho_{2}]=\mathop{\rm tr\,}\nolimits[(D\rho_{1})^{\dagger}\rho_{2}].$
Hence (3.4) is proved.
ii) The calculations above show that the operator $\Delta$} is neither symmetric nor nonpositive. In particular, the calculations (3.1)–(3.3) give in this case
$\mathop{\rm tr\,}\nolimits[\rho D\rho]=|a_{ik}|^{2}[\rho_{i}\rho_{k}-\rho_{k}^{2}]$ (3.5)
which is not enough for the nonpositivity.
4 Injectivity
Let us prove Theorem 2.3 iii). We must show that $\rho=0$ if $D\rho=0$ and $\rho\in{\cal D}$. In the case $D\rho=0$ we have also $\langle\rho,D\rho\rangle=0$. The formula (3.3) gives
$\langle\rho,D\rho\rangle_{\rm HS}=-|a_{ik}|^{2}[\rho_{i}-\rho_{k}]^{2},\qquad\rho\in{\cal D}.$ (4.1)
It implies that $\langle\rho,D\rho\rangle=0$ only if $a_{ik}=0$ for $\rho_{i}\neq\rho_{k}$. Therefore, the eigenspaces corresponding to the eigenvalues $\rho_{i}$ are invariant with respect to $a$ and $a^{\dagger}$. However, there is only one such eigenspace, the entire Hilbert space $X$. Accordingly, all eigenvalues $\rho_{i}$ coinside. Hence, all the eigenvalues are zeros since the zero is in spectrum of any density operator $\rho\in{\cal D}$.
Remark 4.1.
The last step uses the fact that $\rho\in{\cal D}$. Alternatively, we can use the argument that for $\rho\in{\cal D}$ with identical nonzero eigenvalues, the Hilbert–Schmidt norm (2.4) is infinite. This argument also works for general $\rho\in{\rm HS}$ if the relation (4.1) holds. However, our proof of this relation is correct only for $\rho\in{\cal D}$.
5 Acknowledgements
The authors thank S. Kuksin, M.I. Petelin, A. Shnirelman and H. Spohn for longterm fruitful discussions, and the Institute of Mathematics of BOKU University for the support and hospitality. The research is supported by Austrian Science Fund (FWF) PAT 3476224.
6 Conflict of interest
We have no conflict of interest.
7 Data availability statement
The manuscript has no associated data.