### v4

# On dissipation operators of Quantum Optics [^1] The research was carried out within the state assignment of Ministry of Science and Higher Education of the Russian Federation for IITP RAS

Abstract
We consider dissipation operators used in Quantum Optics for the description of quantum spontaneous emission in the context of damped driven Jaynes–Cummings equations. The equations describe quantised one-mode Maxwell field coupled to a two-level molecule.
We nonpositivity of two basic dissipation operators is proved in the framework of the theory of the Hilbert space of the Hilbert-Schmidt operators. We show that one of the operators is symmetric, while the other is not.
MSC classification: 81V80, 81S05, 81S08 37K06, 78A40, 78A60.
Keywords: Quantum Optics; laser; Jaynes–Cummings model; Hamiltonian; density operator; pumping; dissipation; Hilbert space; Hilbert–Schmidt operator; trace.
1 Jaynes–Cummings equation
The Jaynes–Cummings equation (JCE) is basic model of Quantum Optics. The model without damping and pumping was introduced in 1963 by Jaynes and Cummings [6] (the survey can be found in [9]). The damped driven Jaynes–Cummings equations read as [13, (5.107)]:
$\dot{\rho}(t)={\cal A}(t)\rho(t):=-i[H(t),\rho(t)]+\gamma D\rho(t),\qquad t\geq 0.$ (1.1)
[…]
$a|n\rangle=\sqrt{n}|n-1\rangle,\qquad a^{\dagger}|n\rangle=\sqrt{n+1}|n+1\rangle,\qquad[a,a^{\dagger}]=1.$ (1.2)
The Hamiltonian $H(t)$ is the sum
$H(t)\!=\!H_{0}\!+\!pV(t),\,\,\,H_{0}\!=\!H_{F}\!+\!H_{A},\,\,\,H_{F}\!=\!\omega_{c}a^{\dagger}a,\,\,\,H_{A}\!=\!\frac{1}{2}\omega_{a}\sigma_{3},\,\,\,V(t)\!=\!\sigma_{1}(a\!+\!a^{\dagger})\!+\!A^{e}(t).$ (1.3)
[…]
2 Dissipation operators and main results
The dissipation operator $D$ in (1.1) describes the quantum spontaneous emission which provides an energy loss to balance the pumping. The dissipation operators must satisfy the following conditions.
We will consider the dissipation operators [13, (5.107)] and [4, (5.6.20)]:
$D\rho=a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a+a^{\dagger}\rho a-\frac{1}{2}aa^{\dagger}\rho-\frac{1}{2}\rho aa^{\dagger},\qquad{\Delta}\rho=a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a.$ (2.1)
Definition 2.1.
$\rm HS$} is the Hilbert space of Hermitian Hilbert–Schmidt operators with the inner product [11, Thm VI.22 (c)]
$\langle\rho_{1},\rho_{2}\rangle_{\rm HS}=\mathop{\rm tr\,}\nolimits[\rho_{1}\rho_{2}].$ (2.2)
Definition $2.$
[…]
ii) $X_{0}$ is the space of finite linear combinations of the vectors $|n,s_{\pm}\rangle$.
iii) ${\cal D}\subset{\rm HS}$ is the subspace of finite rank Hermitian operators
$\rho=\sum_{n,n^{\prime}}\,\sum_{s,s^{\prime}=s_{\pm}}\rho_{n,s;n^{\prime},s^{\prime}}|n,s\rangle\otimes\langle n^{\prime},s^{\prime}|.$ (2.3)
Every density operator $\rho\in{\rm HS}$ is defined uniquely by its matrix entries
$\rho_{n,s;n^{\prime},s^{\prime}}=\langle n,s|\rho|n^{\prime},s^{\prime}\rangle,$ (2.4)
which are Hermitian $2\times 2$ matrices. The Hilbert–Schmidt norm, corresponding to the inner product (2.2), can be written as
$\|\rho\|_{\rm HS}^{2}=\mathop{\rm tr\,}\nolimits[\rho^{2}]=\sum_{n,n^{\prime}=0}^{\infty}\,\sum_{s,s^{\prime}=s_{\pm}}|\rho_{n,s;n^{\prime},s^{\prime}}|^{2}<\infty.$ (2.5)
The operator $D$ is a modification of the version $\Delta$}. The version $\Delta$} has been introduced in [1][3], and used in [4, (5.6.20)] and [5, (18)], [10], [12, (4)]. The version is symmetric in the following two aspects: for $\rho\in{\cal D}$, ${\Delta}\rho$ is the selfadjoint operator with zero trace that is necessary for the dynamics of $\rho(t)$ in selfadjoint operators with constant trace. The modification $D$ restores the third symmetry with respect to the interchange of $a$ and $a^{\dagger}$.
The main result of present paper is the following theorem.
Theorem 2.3.
i) Both dissipation operators $D$ and $\Delta$} are defined on the domain ${\cal D}\subset{\rm HS}$ and nonpositive:
$\langle\rho,D\rho\rangle_{\rm HS}\leq 0,\qquad\langle\rho,{\Delta}\rho\rangle_{\rm HS}\leq 0,\qquad\rho\in{\cal D}.$ (2.6)
ii) The operator $D:{\cal D}\to{\cal D}$ is symmetric:
$\langle\rho_{1},D\rho_{2}\rangle_{\rm HS}=\langle D\rho_{1},\rho_{2}\rangle_{\rm HS},\qquad\rho_{1},\rho_{2}\in{\cal D},$ (2.7)
while $\Delta$} is not symmetric.
iii) Both operators $D,{\Delta}:{\cal D}\to{\cal D}$ are injective.
Corollary 2.4.
The dissipation operator $D$ admits the selfadjoint Friedrichs extension.
3 Nonpositivity and symmetry
Here we prove Theorem 2.3 i) and ii). First, let us prove the nonpositivity of $\Delta$}. For $\rho\in{\cal D}$,
$\langle\rho,{\Delta}\rho\rangle_{\rm HS}$ $=$ $\mathop{\rm tr\,}\nolimits\big(\rho{\Delta}\rho\big)=\mathop{\rm tr\,}\nolimits\Big(\rho\big(a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a\big)\Big)$ (3.1)
$=$ $\mathop{\rm tr\,}\nolimits\big(\rho a\rho a^{\dagger}-\rho a^{\dagger}a\rho\big)=\mathop{\rm tr\,}\nolimits\big(\rho a\rho a^{\dagger}-a^{\dagger}a\rho^{2}\big).$
Now we use the fact that $\rho$ is a finite rank Hermitian operator (2.3). Then (1.2) implies that the operators $\rho a\rho a^{\dagger}$ and $a^{\dagger}a\rho^{2}$ have only finite number of nonzero entries (2.4), so their traces are well-defined. Moreover, $\rho$ admits a finite spectral resolution in the orthonormal basis of its eigenvectors $e_{i}\in X_{\infty}$:
$\rho=\sum_{i=1}^{\nu}\rho_{i}e_{i}\otimes e_{i}.$ (3.2)
In this basis, the entries $\rho_{ij}=\rho_{i}\delta_{ij}$, and the entries $a_{jk}=\langle e_{j},ae_{k}\rangle$ and $a^{\dagger}_{kl}=\langle e_{k},a^{\dagger}e_{l}\rangle$ of the operators $a$ and $a^{\dagger}$ are well-defined. Hence, (3.1) implies, with summation in repeated indices,
$\langle\rho,{\Delta}\rho\rangle_{\rm HS}$ $=$ $\rho_{i}\delta_{ij}a_{jk}\rho_{k}\delta_{kl}a^{\dagger}_{li}-a^{\dagger}_{kl}a_{lj}\rho_{j}^{2}\delta_{jk}=\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{kl}a_{lk}\rho_{k}^{2}$ (3.3)
$=$ $\rho_{i}a_{ik}\rho_{k}a^{\dagger}_{ki}-a^{\dagger}_{ki}a_{ik}\rho_{k}^{2}=a_{ik}a^{\dagger}_{ki}(\rho_{i}\rho_{k}-\rho_{k}^{2})$
$=$ $\frac{1}{2}\Big(a_{ik}a^{\dagger}_{ki}(\rho_{i}\rho_{k}-\rho_{k}^{2})+a_{ki}a^{\dagger}_{ik}(\rho_{k}\rho_{i}-\rho_{i}^{2}))=-\frac{1}{2}|a_{ik}|^{2}(\rho_{i}-\rho_{k})^{2}\leq 0$
since $a^{\dagger}_{ik}=\overline{a}_{ki}$. Now, the nonpositivity is proved for $\Delta$}.
Lemma 3.1.
$D={\Delta}+{\Delta}^{\dagger}$, where $\Delta$$^{\dagger}$ differs from $\Delta$} by swapping $a$ and $a^{\dagger}$:
${\Delta}^{\dagger}\rho=a^{\dagger}\rho a-\frac{1}{2}\rho a^{{\dagger}}a-\frac{1}{2}a^{{\dagger}}a\rho,\qquad\rho\in{\cal D}.$ (3.4)
Proof.
For $\rho_{1},\rho_{2}\in{\cal D}$,
$\mathop{\rm tr\,}\nolimits[\rho_{1}({\Delta}\rho_{2})]$ $=$ $\mathop{\rm tr\,}\nolimits\Big(\rho_{1}(a\rho_{2}a^{\dagger}-\frac{1}{2}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{2}a^{{\dagger}}a)\Big)$
$=$ $\mathop{\rm tr\,}\nolimits\Big(\rho_{1}a\rho_{2}a^{\dagger}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}\rho_{1}\rho_{2}a^{{\dagger}}a\Big)$
$=$ $\mathop{\rm tr\,}\nolimits\Big(a^{\dagger}\rho_{1}a\rho_{2}-\frac{1}{2}\rho_{1}a^{{\dagger}}a\rho_{2}-\frac{1}{2}a^{{\dagger}}a\rho_{1}\rho_{2}\Big)$
$=$ $\mathop{\rm tr\,}\nolimits\Big((a^{\dagger}\rho_{1}a-\frac{1}{2}\rho_{1}a^{{\dagger}}a-\frac{1}{2}a^{{\dagger}}a\rho_{1})\rho_{2}\Big)=\mathop{\rm tr\,}\nolimits[({\Delta}^{\dagger}\rho_{1})\rho_{2}].$
Hence, (3.4) is proved. ∎
As the corollary, the nonpositivity of $\Delta$$^{\dagger}$ holds by the same arguments as for $\Delta$} just by swapping $a$ and $a^{\dagger}$, and hence, $D={\Delta}+{\Delta}^{\dagger}$ is also nonpositive. Moreover, $D$ is symmetric on $\cal D$}, while $\Delta$} is not.
Remark 3.2.
The proof of the nonpositivity essentially depends on the symmetry of $\rho$.
4 Injectivity
Let us prove Theorem 2.3 iii). The formula (3.3) gives
$\langle\rho,{\Delta}\rho\rangle_{\rm HS}=-\frac{1}{2}|a_{ik}|^{2}[\rho_{i}-\rho_{k}]^{2},\qquad\rho\in{\cal D}.$ (4.1)
It implies that $\langle\rho,{\Delta}\rho\rangle=0$ only if $a_{ik}=0$ for $\rho_{i}\neq\rho_{k}$. Therefore, the eigenspaces of $\rho$ corresponding to the eigenvalues $\rho_{i}$ are invariant with respect to $a$ and $a^{\dagger}$. However, there is only one such eigenspace, the entire Hilbert space $X$. Accordingly, all eigenvalues $\rho_{i}$ coinside. Hence, all the eigenvalues are zeros since the zero is in spectrum of any density operator $\rho\in{\cal D}$.
Remark 4.1.
The last step uses the fact that $\rho\in{\cal D}$. Alternatively, we can use the argument that the Hilbert–Schmidt norm (2.5) is infinite if all the eigenvalues $\rho_{i}$ are nonzero and identical. This argument also works for general $\rho\in{\rm HS}$ if the relation (4.1) holds. However, our proof of this relation is correct only for $\rho\in{\cal D}$.
5 Acknowledgements
The authors thank S. Kuksin, M.I. Petelin, A. Shnirelman and H. Spohn for longterm fruitful discussions.